OpenStudy (dtan5457):

How do I and graph the function of y=-5/x and compare it to y=1/x?

OpenStudy (dtan5457):

@DanJS

OpenStudy (dtan5457):

@SolomonZelman @Jhannybean

OpenStudy (dtan5457):

I mean i know how to graph it but im not sure how am i suppose to compare these two?

OpenStudy (jhannybean):

First graph 1/x

OpenStudy (ahsome):

Simply graph and see how they look different. Just from looking at this, I know that one is goinf to be steeper (-5/x) and one is going to be positive direction (Going Up, which is 1/x) and one is going to be negative direction (Goind Down, which is -5/x)

OpenStudy (ahsome):

 

OpenStudy (ahsome):

Simply state the differences, what direction they are going etc.

OpenStudy (dtan5457):

anything to do asymptotes or are they the same?

OpenStudy (ahsome):

You coudl state where the asymtopes are, if there are that are vertical, horizontal or oblique

OpenStudy (solomonzelman):

will give it a notation: f(x)=1/x g(x)=-5/x you multiplied the f(x) times 5, (stretch), then reflected it across the x axis (by multiplying it times -1). Boom, there is your g(x)...

OpenStudy (dtan5457):

to clarify, if its bigger than 1 as a in this case, it will be stretched, less than 1=shrunk?